Two charges A and B are fixed in place, at different distances from a certain spot. At this spot the potentials due to the two charges are equal. Charge A is 0.162 m from the spot, while charge B is 0.406 m from it. Calculate the ratio qB/qA of the charges.
Answer #1

we know, Potential due to two charges = kqA/rA + kqB/rB =0

  kqA/0.162 + kqB/0.406 = 0

kqA/0.162 = - kqB/0.406

qB/qA = -0.406/0.162 = -2.506 Answer